OpenStudy (anonymous):

Solve for n: 6(5n – 1) = 2(15n – 3)

OpenStudy (anonymous):

(30n-6)= (30n-6)...they're both equal so wouldn't n=1

OpenStudy (anonymous):

6(5n – 1) = 2(15n – 3) 30n-6=30n-6 30n-30n=-6+6 0=0 ^^

OpenStudy (anonymous):

-1 or 3 or no solution or all real solutions are the anwsers

OpenStudy (anonymous):

no sol

OpenStudy (anonymous):

lol